For a person at rest, the blood pressure in the aorta is 120 torr, and the velocity of the blood is 25cm/s. The person then goes for the brisk walk raising the blood velocity to 100cm/s. Using Bernoulli's equation and assuming that the height of the blood is constant calculate the blood pressure in the aorta during the brisk walk. Hint: convert everything to standard units.

1
Expert's answer
2021-11-24T12:01:51-0500

"120\ torr=15998.68\ Pa"


"p_1+\rho v_1^2/2=p_2+\rho v_2^2/2\to"


"p_2=p_1+\rho v_1^2/2-\rho v_2^2/2=p_1+(\rho/2)(v_1^2-v_2^2)="


"=15998.68+(1060/2)(0.25^2-1^2)=15501.8\ (Pa)" . Answer